2x2 = 3y2 + 5 , x - 3 + 5 = 0
Answer by KnightOwlTutor(293)

x-3+5=0 simplify to get x+2=0 subtract 2 from both sides and get x=-2
replace each x with -2. When you square a -number the result is positive.
2(-2)^2=3y^2+5 8=3y^2+5
Subtract 5 from both side and you get 3=3y^2
Divide both sides by 3
1=y^2 y =1 or -1.